A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is
55%
salt and Solution B is
80%
salt. She wants to obtain
160
ounces of a mixture that is
60%
salt. How many ounces of each solution should she use?

Found 3 solutions by ikleyn, adamhen894, Alan3354:
Answer by ikleyn(52787)

.
It is useful to know that
    The maximum solubility of table salt (sodium chloride) in water at 20 degrees Celsius is 35.7 grams per 100 milliliters of water. 

https://www.reference.com/geography/maximum-solubility-salt-water-b5cba303939f22e


So, at normal conditions the maximum salt solubility is about 35 - 36 percents (by mass).


------------------
There is entire bunch of lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for alloys
    - Typical word problems on mixtures from the archive
in this site.

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook in the section "Word problems" under the topic "Mixture problems".


Answer by adamhen894(15)

solutionA
x ounces of 55%salt
solutionB
(160-x)ounces of 80% salt
mixture of two solution
160 ounces of 60% salt
now we can set up the equation, solve for x
0.55x+0.8(160-x)=0.6*160
x= 128 ounces of 55% salt
and 32 ounces of 60% salt
input x value into equation, they equal to each other.
if you have alternative way of solving it, i would love to hear
adamchen894@gmail.com

Answer by Alan3354(69443)

A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is
55% salt and Solution B is 80% salt.
She wants to obtain 160 ounces of a mixture that is
60% salt. How many ounces of each solution should she use?
==============
Ignoring the fact that 80% salt and 55% salt is not possible:
---
e = amount of 80%
f = amount of 55%
---
e + f = 160 (total solution)
80e + 55f = 60*160 (total salt)
-----
16e + 11f = 1920
f = 160 - e
16e + 11*(160-e) = 1920
16e + 1760 - 11e = 1920
5e = 160
e = 32 ounces
f = 128 ounces